OpenStudy (anonymous):

Which ordered pair is a solution of the equation: y-3=x+2 a. (5, 10) b. (0, 0) c. (7, 6) d. (10, 5)

rebeccaxhawaii (rebeccaxhawaii):

plug it in for x and y and if its a true statement then its a solution

OpenStudy (xguardians):

You have to substitute y-3=x+2 Add 3 to both sides y = x+5 Plug in values that are 5 apart (5,10) and (10,5) Try A) y-3 = x+2 10 - 3 = 5 + 2 7=7 A is the answer.